Ken drove his car from home to Chicago at a rate of 40mph and returned at a rate of 45 mph. If his time going exceeded his time returning by 30 mins find his time going and returning?
Answer by oscargut(686)

Ken drove his car from home to Chicago at a rate of 40mph and returned at a rate of 45 mph. If his time going exceeded his time returning by 30 mins find his time going and returning?
let t1 time on going and t2 time on returning
the distance is the same going and returning so

40t1=45t2
t1=t2+0.5
then 40(t2+0.5)=45t2 then -5t2+20=0 then t2=4 and t1=4.5